If Qs = -20 + 10p, and Qd = 400 - 20p, what is the equilibrium price and quantity?


1
Expert's answer
2021-06-13T17:39:01-0400

At the equilibrium:

"Q_s=Q_d \\

-20+10p=400-20p \\

30p=420 \\

p_{eq}=14 \\

Q_{eq} = -20+ 10 \times 14 = 120"